# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 5.1. Intrinsic mirror symmetry for projective degenerations
The intrinsic mirror symmetry construction particularly applies to a projective morphism $g\colon\mathcal{X}\to S$ from a smooth variety $\mathcal{X}$ with $K_{\mathcal{X}}=0$ to a non-singular, non-complete curve $S$ with single closed point $0\in S$, such that the central fiber $\mathcal{X}_{0}=g^{-1}(0)$ is a reduced simple normal crossing divisor on $\mathcal{X}$ – this is a special situation of a so called “log smooth” morphism as in [47, §1.1, pg 7]. In particular, the schemes $\mathcal{X}$ and $S$ are viewed as “log schemes”, with a log structure defined by the data of the divisors $\mathcal{X}_{0}\subset\mathcal{X}$ and $0\in S$ respectively.
@@ 5.2. Mirror algebras
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. As $\pi$ is a smoothing of $\mathcal{X}_{0}$, there is a natural log structure on $\mathcal{X}_{0}$, log smooth over the standard log point – see e.g. [92, §3.2] for foundational aspects of log structures on formal schemes. We denote by $\mathcal{X}_{0}^{\dagger}$ the corresponding log scheme. Its tropicalization $\Sigma(\mathcal{X}_{0}^{\dagger})$ is the cone $(C(P),C(\mathscr{P}))$ over the dual intersection complex $(P,\mathscr{P})$.
@@
The theory of punctured maps is developed in [2] under the assumption that the target has a log structure defined with respect to the Zariski topology (and not with respect to the étale topology). Correspondingly, the mirror construction of Gross–Siebert [47, 48] is formulated under the assumption that the central fiber is simple normal crossing instead of just normal crossing. If the open Kulikov surface $\mathcal{X}_{0}$ is not simple normal crossing, there is an edge $e$ in $\mathscr{P}$ connecting a vertex $v$ to itself. Adding a ray passing through a rational point in the interior of every such an edge $e$ defines a log blow-up $\widetilde{\mathcal{X}}$ of $\mathcal{X}$, so that the central fiber $\widetilde{\mathcal{X}}_{0}$ is now simple normal crossing, and the Gross–Siebert construction (with the modifications described below to deal with the non-compactness of $\mathcal{X}_{0}$) applies to such a degeneration $\widetilde{\mathcal{X}}\rightarrow\Delta$. Since this log blow-up is harmless for all our arguments, we will no longer mention it explicitly for the remaining of this paper.
@@
Let $r\in C(P)_{\mathbb{Z}}\setminus\{0\}$, and denote by $Z_{r}$ the stratum of $\mathcal{X}_{0}^{\dagger}$ corresponding to the smallest cone of $C(P)$ containing $r$. As in [47, §3.1], one can define a moduli space $\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)$ of punctures in $\mathcal{X}_{0}$ with contact order $-r$, which is a log stack with underlying stack $Z_{r}\times B\mathbb{G}_{m}$. As in Definition [47, 3.5], for every type $\tau$ of genus zero containing a punctured point $x_{\mathrm{out}}$ with contact order $-r$. there is a log evaluation morphism at $x_{\mathrm{out}}$
@@
By [47, Proposition 3.8], a choice of a point $z$ in the interior of $Z_{r}$ induces a log morphism
@@
Let $\nu\colon C\to\mathcal{X}_{0}^{\dagger}$ be a punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$. As the class $\beta\in NE(\mathcal{X}_{0})$ of the image $\nu(C)$ is contained in $NE(\mathcal{X}_{0}/\overline{\mathcal{X}}_{0})$, it follows that $\nu(C)$ is contracted to a point under the map $f\colon\mathcal{X}\to\overline{\mathcal{X}}$. By definition of the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$, the image $\nu(C)$ contains the point $z$. Thus, we have $f(\nu(C))=\{f(z)\}$, and so $\nu(C)\subset f^{-1}(\{f(z)\})$. Note that for $r\neq 0$ we have $Z_{r}\subset\mathcal{X}_{0}$ and since $z\in Z_{r}$ we have $z\in\mathcal{X}_{0}$. Therefore, we have $f(z)\in\overline{\mathcal{X}}_{0}$, and so $f^{-1}(\{f(z)\})$ is actually a subvariety of $\mathcal{X}_{0}$. Since every punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ factors through the subvariety $V:=f^{-1}(\{f(z)\})$ of $\mathcal{X}_{0}$, it follows that
$\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)\simeq\mathcal{M}(V^{\dagger},\tau,z)\,,$
where $V^{\dagger}$ is the log scheme obtained by restricting the log structure of $\mathcal{X}_{0}^{\dagger}$ to $V=f^{-1}(\{f(z)\})$. Since $f\colon\mathcal{X}\to\overline{\mathcal{X}}$ is projective, the fiber $V=f^{-1}(\{f(z)\})$ of $f$ is a projective subvariety of $\mathcal{X}_{0}$, and so the properness of $\mathcal{M}(V^{\dagger},\tau,z)$ follows by [2]. ∎
@@
All the arguments in the proof of [47, Theorem 1.9] for projective degenerations $g:\mathcal{X}\rightarrow S$ as in §5.1 can be adapted to the set-up of open Kulikov degenerations to prove Theorem 5.5. Indeed, the proof of [47, Theorem 1.9] for a projective degeneration $g:\mathcal{X}\rightarrow S$ is based on the study of the moduli spaces $\mathcal{M}(\mathcal{X},\tau,z)$, where $\tau$ is a genus zero type of punctured maps with three marked points and an additional punctured point. For an open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$, the corresponding moduli spaces $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ are proper by Theorem 5.2 and so the non-compactness of $\mathcal{X}_{0}$ is not an issue. ∎
@@ 5.3. Structure constants from the canonical scattering diagram
where $n_{W}$ is the primitive normal vector to $C(W)$, positive on the half-space containing $b_{j}$. Finally, we denote by $b_{\mathrm{out}}$ the final domain of linearity of $b$, and $c_{b_{\mathrm{out}}}z^{m_{b_{\mathrm{out}}}}$ the corresponding monomial. By [48, Theorem 6.1], for every $p,q,r\in C(P)_{\mathbb{Z}}$, we then have
@@
As reviewed above, by [48, Theorem 6.1], $N_{pqr}^{\beta}(\mathcal{X})$ can be computed from the canonical scattering diagram, which only depends on $\mathcal{X}_{0}^{\dagger}$ by [48, Proposition 3.7]. ∎
@@
Lemma 5.6 does not immediately follows from Definition 5.3. Indeed, even if $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ only depends on $\mathcal{X}_{0}^{\dagger}$, the definition of the relative obstruction theory and so of the virtual fundamental class $[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)]^{\mathrm{virt}}$ a priori depends on $\mathcal{X}$.
@@ 6. Polarized mirrors of open Kulikov degenerations
After classifying in §6.1 the elementary flops between quasi-projective open Kulikov degenerations, we study in §6.2 how the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ of the mirror algebra change under these flops. This is used in §6.3 to prove a key finiteness property: the mirror algebra of a quasi-projective open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ can be defined over $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$, and not just over the Artinian rings $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]/I$. In §6.4-6.5, we define the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$ as a coarsening of the Mori fan $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$, and we show that the Proj of the mirror algebras of the various projective crepant resolutions of $\overline{\mathcal{X}}$ glue and extend to a polarized mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{sec}})\rightarrow S_{\mathcal{X}}^{\mathrm{sec}}$ over the complete toric variety $S_{\mathcal{X}}^{\mathrm{sec}}$ with fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$.
@@ 6.1. Elementary flops
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, and let $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be an elementary small $\mathbb{Q}$-factorial modification of $\mathcal{X}$ over $\overline{\mathcal{X}}$. Then, the exceptional locus of $b$ is a smooth rational curve $C\subset\mathcal{X}_{0}$, such that
@@
    either $C$ is an internal exceptional curve of an irreducible component of $\mathcal{X}_{0}$,
@@
    or $C$ is an irreducible component $C=V_{i}\cap V_{j}$ of the double locus of $\mathcal{X}_{0}$, intersection of two irreducible components $V_{i}$, $V_{j}$ of $\mathcal{X}_{0}$, and such that $\deg N_{C|V_{i}}=\deg N_{C|V_{j}}=-1$.
@@
As $\mathcal{X}_{0}$ is generic as in Definition 3.19, and so does not contain any internal $(-2)$-curve, it follows from the proof of [85, Lemma 3.2] that the connected components of the exceptional locus of $b$ are either as in i) or ii). Hence, to prove Lemma 6.1, it is enough to show that the exceptional locus of $b$ is in fact connected. Assume that $C$ and $C^{\prime}$ are two curves as in i) or ii) with proportional classes. Curves as in i) have negative intersection with exactly one irreducible component of $\mathcal{X}_{0}$, whereas curves as in ii) have negative intersection with exactly two irreducible components of $\mathcal{X}_{0}$. Hence, either both $C$ and $C^{\prime}$ are as in i), or both $C$ and $C^{\prime}$ are as in ii). In the latter case, we have $C=C^{\prime}=V_{i}\cap V_{j}$, where $V_{i}$ and $V_{j}$ are the two irreducible components of $\mathcal{X}_{0}$ with $C\cdot V_{i}<0$ and $C\cdot V_{j}<0$. In the former case, both $C$ and $C^{\prime}$ are internal exceptional curves in the same irreducible component $V_{i}$ of $\mathcal{X}_{0}$, the one such that $C\cdot V_{i}<0$, and intersect the same irreducible component $V_{i}\cap V_{j}$ of $\partial V_{i}$: $V_{j}$ is the unique irreducible component of $\mathcal{X}_{0}$ such that $C.V_{i}>0$. Hence, the line bundle $\mathcal{O}_{V_{i}}(C-C^{\prime})$ is trivial in restriction to every irreducible component of $\partial V_{i}$, and so, by the description of $\mathrm{Pic}(\mathcal{X}_{0})$ given by Lemma 3.10, defines a line bundle on $\mathcal{X}_{0}$. By Lemma 3.10, we also have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so this line bundle formally deforms into a line bundle $L$ on $\mathcal{X}$. If $C\neq C^{\prime}$, then $L\cdot C=-1$, whereas $L\cdot C^{\prime}=1$, and so, as the classes $C$ and $C^{\prime}$ are assumed to be proportional, $C=-C^{\prime}$, in contradiction with the strict convexity of the cone of curves $NE(\mathcal{X}/\overline{\mathcal{X}})$. ∎
Following [7, Definition 7.15], we call M1 flop a small $\mathbb{Q}$-factorial modification as in Lemma 6.1 i), that is the flop of an internal exceptional curve, and M2 flop a small $\mathbb{Q}$-factorial modification as in Lemma 6.1 ii), that is the flop of an exceptional curve contained in the double locus. In restriction to the central fiber $\mathcal{X}_{0}$, M1 and M2 flops act like the M1 and M2 modifications of open Kulikov surfaces introduced in §3.3. Note that M1 and M2 flops are often referred to as Type 1 and Type 2 modifications in the literature [86, pg. 12].

### 30 Mar 2024 v2

@@ 5.1. Intrinsic mirror symmetry for projective degenerations
The intrinsic mirror symmetry construction particularly applies to a projective morphism $g\colon\mathcal{X}\to S$ from a smooth variety $\mathcal{X}$ with $K_{\mathcal{X}}=0$ to a non-singular, non-complete curve $S$ with single closed point $0\in S$, such that the central fiber $\mathcal{X}_{0}=g^{-1}(0)$ is a reduced simple normal crossing divisor on $\mathcal{X}$ – this is a special situation of a so called “log smooth” morphism as in [48, §1.1, pg 7]. In particular, the schemes $\mathcal{X}$ and $S$ are viewed as “log schemes”, with a log structure defined by the data of the divisors $\mathcal{X}_{0}\subset\mathcal{X}$ and $0\in S$ respectively.
@@ 5.2. Mirror algebras
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. As $\pi$ is a smoothing of $\mathcal{X}_{0}$, there is a natural log structure on $\mathcal{X}_{0}$, log smooth over the standard log point – see e.g. [93, §3.2] for foundational aspects of log structures on formal schemes. We denote by $\mathcal{X}_{0}^{\dagger}$ the corresponding log scheme. Its tropicalization $\Sigma(\mathcal{X}_{0}^{\dagger})$ is the cone $(C(P),C(\mathscr{P}))$ over the dual intersection complex $(P,\mathscr{P})$.
@@
The theory of punctured maps is developed in [2] under the assumption that the target has a log structure defined with respect to the Zariski topology (and not with respect to the étale topology). Correspondingly, the mirror construction of Gross–Siebert [48, 49] is formulated under the assumption that the central fiber is simple normal crossing instead of just normal crossing. If the open Kulikov surface $\mathcal{X}_{0}$ is not simple normal crossing, there is an edge $e$ in $\mathscr{P}$ connecting a vertex $v$ to itself. Taking the cone $C(\widetilde{\mathscr{P}})$ over a rational triangulation $\widetilde{\mathscr{P}}$ refining $\mathscr{P}$ and containing as vertices some rational points in the interior of every such edges $e$ defines a log blow-up $\widetilde{\mathcal{X}}$ of $\mathcal{X}$, so that the central fiber $\widetilde{\mathcal{X}}_{0}$ is now simple normal crossing, and the Gross–Siebert construction (with the modifications described below to deal with the non-compactness of $\mathcal{X}_{0}$) applies to such a degeneration $\widetilde{\mathcal{X}}\rightarrow\Delta$. Since this log blow-up is harmless for all our arguments, we will no longer mention it explicitly for the remaining of this paper.
@@
Let $r\in C(P)_{\mathbb{Z}}\setminus\{0\}$, and denote by $Z_{r}$ the stratum of $\mathcal{X}_{0}^{\dagger}$ corresponding to the smallest cone of $C(P)$ containing $r$. As in [48, §3.1], one can define a moduli space $\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)$ of punctures in $\mathcal{X}_{0}$ with contact order $-r$, which is a log stack with underlying stack $Z_{r}\times B\mathbb{G}_{m}$. As in Definition [48, 3.5], for every type $\tau$ of genus zero containing a punctured point $x_{\mathrm{out}}$ with contact order $-r$. there is a log evaluation morphism at $x_{\mathrm{out}}$
@@
By [48, Proposition 3.8], a choice of a point $z$ in the interior of $Z_{r}$ induces a log morphism
@@
Let $\nu\colon C\to\mathcal{X}_{0}^{\dagger}$ be a punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$. As the class $\beta\in NE(\mathcal{X}_{0})$ of the image $\nu(C)$ is contained in $NE(\mathcal{X}_{0}/\overline{\mathcal{X}}_{0})$, it follows that $\nu(C)$ is contracted to a point under the map $f\colon\mathcal{X}\to\overline{\mathcal{X}}$. By definition of the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$, the image $\nu(C)$ contains the point $z$. Thus, we have $f(\nu(C))=\{f(z)\}$, and so $\nu(C)\subset f^{-1}(\{f(z)\})$. Note that for $r\neq 0$ we have $Z_{r}\subset\mathcal{X}_{0}$ and since $z\in Z_{r}$ we have $z\in\mathcal{X}_{0}$. Therefore, we have $f(z)\in\overline{\mathcal{X}}_{0}$, and so $f^{-1}(\{f(z)\})$ is actually a subvariety of $\mathcal{X}_{0}$. Since every punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ factors through the subvariety $V:=f^{-1}(\{f(z)\})$ of $\mathcal{X}_{0}$, it follows that $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)\simeq\mathcal{M}(V^{\dagger},\tau,z)$, where $V^{\dagger}$ is the log scheme obtained by restricting the log structure of $\mathcal{X}_{0}^{\dagger}$ to $V=f^{-1}(\{f(z)\})$. Since $f\colon\mathcal{X}\to\overline{\mathcal{X}}$ is projective, the fiber $V=f^{-1}(\{f(z)\})$ of $f$ is a projective subvariety of $\mathcal{X}_{0}$, and so the properness of $\mathcal{M}(V^{\dagger},\tau,z)$ follows by [2]. ∎
@@
All the arguments in the proof of [48, Theorem 1.9] for projective degenerations $g:\mathcal{X}\rightarrow S$ as in §5.1 can be adapted to the set-up of open Kulikov degenerations to prove Theorem 5.5. Indeed, the proof of [48, Theorem 1.9] for a projective degeneration $g:\mathcal{X}\rightarrow S$ is based on the study of the moduli spaces $\mathcal{M}(\mathcal{X},\tau,z)$, where $\tau$ is a genus zero type of punctured maps with three marked points and an additional punctured point. For an open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$, the corresponding moduli spaces $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ are proper by Theorem 5.2 and so the non-compactness of $\mathcal{X}_{0}$ is not an issue. ∎
@@ 5.3. Structure constants from the canonical scattering diagram
where $n_{W}$ is the primitive normal vector to $C(W)$, positive on the half-space containing $b_{j}$. Finally, we denote by $b_{\mathrm{out}}$ the final domain of linearity of $b$, and $c_{b_{\mathrm{out}}}z^{m_{b_{\mathrm{out}}}}$ the corresponding monomial. By [49, Theorem 6.1], for every $p,q,r\in C(P)_{\mathbb{Z}}$, we then have
@@
As reviewed above, by [49, Theorem 6.1], $N_{pqr}^{\beta}(\mathcal{X})$ can be computed from the canonical scattering diagram, which only depends on $\mathcal{X}_{0}^{\dagger}$ by [49, Proposition 3.7]. ∎
@@
Lemma 5.6 does not immediately follows from Definition 5.3. Indeed, even if $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ only depends on $\mathcal{X}_{0}^{\dagger}$, the definition of the relative obstruction theory, and so also the definition of the virtual fundamental class $[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)]^{\mathrm{virt}}$, depend a priori on $\mathcal{X}$.
@@ 6. Polarized mirrors of open Kulikov degenerations
After classifying in §6.1 the elementary flops between quasi-projective open Kulikov degenerations, we study in §6.2 how the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ of the mirror algebra change under these flops. This is used in §6.3 to prove a key finiteness property: the mirror algebra of a quasi-projective open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ can be defined over $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$, and not just over the Artinian rings $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]/I$. In §6.4-6.5, we define the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$ as a coarsening of the Mori fan $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$. We show that there is a natural stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ with underlying fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, and that the Proj of the mirror algebras of the various projective crepant resolutions of $\overline{\mathcal{X}}$ glue and extend to a polarized mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{sec}})\rightarrow\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ over the toric Deligne–Mumford stack $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ with fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$.
@@ 6.1. Elementary flops
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, and let $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be an elementary small $\mathbb{Q}$-factorial modification of $\mathcal{X}$ over $\overline{\mathcal{X}}$. Then, the exceptional locus of $b$ is a smooth rational curve $C\subset\mathcal{X}_{0}$, such that either
@@
    $C$ is an internal exceptional curve of an irreducible component of $\mathcal{X}_{0}$, or
@@
    $C=V_{i}\cap V_{j}$ is an irreducible component of the double locus of $\mathcal{X}_{0}$, intersection of two irreducible components $V_{i}$, $V_{j}$ of $\mathcal{X}_{0}$, and such that $\deg N_{C|V_{i}}=\deg N_{C|V_{j}}=-1$.
@@
As $\mathcal{X}_{0}$ is generic as in Definition 3.19, and so does not contain any internal $(-2)$-curve, it follows from the proof of [86, Lemma 3.2] that the connected components of the exceptional locus of $b$ are either as in i) or ii). Hence, to prove Lemma 6.1, it is enough to show that the exceptional locus of $b$ is in fact connected. Assume that $C$ and $C^{\prime}$ are two curves as in i) or ii) with proportional classes. Curves as in i) have negative intersection with exactly one irreducible component of $\mathcal{X}_{0}$, whereas curves as in ii) have negative intersection with exactly two irreducible components of $\mathcal{X}_{0}$. Hence, either both $C$ and $C^{\prime}$ are as in i), or both $C$ and $C^{\prime}$ are as in ii). In the latter case, we have $C=C^{\prime}=V_{i}\cap V_{j}$, where $V_{i}$ and $V_{j}$ are the two irreducible components of $\mathcal{X}_{0}$ with $C\cdot V_{i}<0$ and $C\cdot V_{j}<0$. In the former case, both $C$ and $C^{\prime}$ are internal exceptional curves in the same irreducible component $V_{i}$ of $\mathcal{X}_{0}$, the one such that $C\cdot V_{i}<0$, and intersect the same irreducible component $V_{i}\cap V_{j}$ of $\partial V_{i}$: $V_{j}$ is the unique irreducible component of $\mathcal{X}_{0}$ such that $C.V_{i}>0$. Hence, the line bundle $\mathcal{O}_{V_{i}}(C-C^{\prime})$ is trivial in restriction to every irreducible component of $\partial V_{i}$, and so, by the description of $\mathrm{Pic}(\mathcal{X}_{0})$ given by Lemma 3.10, defines a line bundle on $\mathcal{X}_{0}$. By Lemma 3.10, we also have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so this line bundle formally deforms into a line bundle $L$ on $\mathcal{X}$. If $C\neq C^{\prime}$, then $L\cdot C=-1$, whereas $L\cdot C^{\prime}=1$, and so, as the classes $C$ and $C^{\prime}$ are assumed to be proportional, $C=-C^{\prime}$, in contradiction with the strict convexity of the cone of curves $NE(\mathcal{X}/\overline{\mathcal{X}})$. ∎
Following [7, Definition 7.15], we call M1 flop a small $\mathbb{Q}$-factorial modification as in Lemma 6.1 i), that is the flop of an internal exceptional curve, and M2 flop a small $\mathbb{Q}$-factorial modification as in Lemma 6.1 ii), that is the flop of an exceptional curve contained in the double locus. In restriction to the central fiber $\mathcal{X}_{0}$, M1 and M2 flops act like the M1 and M2 modifications of open Kulikov surfaces introduced in §3.3. Note that M1 and M2 flops are often referred to as Type 1 and Type 2 modifications in the literature [87, pg. 12].